OpenStudy (ashleyisakitty):

There were 144 tickets sold for a movie. The ratio of the number of child tickets to the number of adult tickets to the number of senior tickets was 4 : 7 : 1. How many of each type of ticket were sold?

OpenStudy (raden):

to child tickets = 4/(4+7+1) * 144 = 4/12 * 144 = .. to adult tickets = 7/(4+7+1) * 144 = 7/12 * 144 = .. to senior tickets = 1/(4+7+1) * 144 = 1/12 * 144 = ..

OpenStudy (whpalmer4):

We know that C + A + S = 144 (child, adult, senior tickets respectively). We divide the tickets into 12 parts (because 4 + 7 + 1 = 12) and that gives us 4/12 children's tickets, 7/12 adult tickets and 1/12 senior tickets. Multiply each fraction by 144 and you'll have the number of tickets sold in each type.